OpenStudy (anonymous):

how do you put Y+3=-5/6(x-1) into standard form?

OpenStudy (anonymous):

do you know what the standard form is? I would usually say it is y = mx + b, where m is the slope and b is the y intercept. However, if your class has said something different, we can use that...

OpenStudy (anonymous):

they are using y=mx+b

OpenStudy (anonymous):

good :) So you need to multiply out and simplify the equation they gave you: Y+3=-5/6(x-1) <<-- is the (x-1) in the denominator or numerator??

OpenStudy (anonymous):

so...i got Y+3=-5/6X+5/6 WHATS THE NEXT STEP?

OpenStudy (anonymous):

guess it was the numerator. So then subtract 3 from both sides to get y all alone on the left. y +3 -3 = -(5/6)x +(5/6) - 3 y = -(5/6)x +(5/6) - 3 Then get a common denominator and combine (5/6) and -3 and you will be done

OpenStudy (anonymous):

Y=-5/6X-13/6.... AND THEN IM STUCK?

OpenStudy (anonymous):

no, you are done :) That is the standard form: y = -(5/6)x - (13/6) is of the same form as... y = mx + b where m = -5/6 and b = -(13/6) That's all there is... you found the standard form :)

OpenStudy (anonymous):

THATS ACTUALLY POINT SLOPE FORM. STANDARD FORM IS Ax+By=c and I dont know how to solve for that

OpenStudy (anonymous):

sorry.:(

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

ok... you are right, it is Ax+By=C no fractions, and A has to be positive.

OpenStudy (anonymous):

I would clear the fractions first so that we just don't have to worry about that any more. Clear your fractions by multiplying by 6 on both sides.

OpenStudy (anonymous):

I get\[6y+18=-5(x-1)\]Are you with me?

OpenStudy (anonymous):

y=-5x-13?

OpenStudy (anonymous):

I'm starting from what you wrote in your original post. y+3=(-5/6)(x-1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is point-slope form. We need to get every variable over to the left with no fractions, in alphabetical order ( x then y ) and with no negative in front of the x term.

OpenStudy (anonymous):

I want to get rid of the fraction first, so we can just forget about fractions.

OpenStudy (anonymous):

ok but how did u get rid of the fractions on the right side

OpenStudy (anonymous):

If the fraction involved has a 6 in the denominator, I can kill it by multiplying both sides by 6.

OpenStudy (anonymous):

This uses the "balanced equation" concept to keep both sides equal... I multiply the right by 6 to kill the fraction. I multiply the left side by 6 to keep things balanced.

OpenStudy (anonymous):

ok got it an then..

OpenStudy (anonymous):

i get 6y+18=-5x+1

OpenStudy (anonymous):

So the result should be 6y+18=(-5)(x-1)

OpenStudy (anonymous):

You need to distribute that -5 to both terms inside the parenthesis.

OpenStudy (anonymous):

6y+18=-5x+5

OpenStudy (anonymous):

so for my answer i get 5x+6y=-19 is this right?

OpenStudy (anonymous):

not quite...

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

you have the x and y parts right

OpenStudy (anonymous):

but if you distribute correctly... the number to use is... subtract 5 from both sides.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

???

OpenStudy (anonymous):

I mean subtract 18 from both sides... der

OpenStudy (anonymous):

so 5x+6y=-17

OpenStudy (anonymous):

check that distribution step... 6y+18=-5(x-1) leads to 6y+18=-5x+5

OpenStudy (anonymous):

then move the x term 5x+6y+18=5

OpenStudy (anonymous):

then subtract 18 from both sides 5x+6y=5-18

OpenStudy (anonymous):

5-18=?

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so... 5x+6y=-13

OpenStudy (anonymous):

that is standard form

OpenStudy (anonymous):

thank you soo much:)

OpenStudy (anonymous):

You're welcome! :)